OpenStudy (anonymous):

The number of traffic accidents that occur on a particular stretch of road during a month follows a Poisson distribution with a mean of 7.6. Find the probability that less than three accidents on this stretch of road next month.

OpenStudy (mathmate):

lambda=7.6 P(X<3)=P(X=0)+P(X=1)+P(X=2) For the poisson distribution, P(X=x)=e^(-lambda)*lambda^(x)/x! or you can look up tables for lambda=7.6 and x=0,1,2 respectively (0.0005, 0.0038, 0.0145)